Mike's age, decreased by the age of his 4 year-old sister, is 11. Mike's age, increased by his uncle's age, is 53. What age is mike's uncle?
Answer by TimothyLamb(4379)

x = uncle age
y = mike age
---
y - 4 = 11
y + x = 53
---
put the system of linear equations into standard form
---
y = 15
x + y = 53
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = uncle age = 38
y = mike age = 15
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---